OpenStudy (anonymous):

how to find the function of the set 1, 6, 31, 156, 781... Please show your work, thanks

OpenStudy (asnaseer):

see if you can spot a pattern between the differences between each successive term of this sequence.

OpenStudy (anonymous):

yes, but HOW to find, for example, the 20th number on the list

OpenStudy (asnaseer):

so what pattern did you spot?

OpenStudy (anonymous):

previous numer x 5 + 1 = next number

OpenStudy (asnaseer):

yes, so you have:\[a_n=5a_{n-1}+1\]with \(a_0=0\)

OpenStudy (asnaseer):

there is a technique to turn this recursive formula into an absolute formula

OpenStudy (asnaseer):

I am not an expert in this but it is described here: http://hcmop.wordpress.com/2012/04/20/using-characteristic-equation-to-solve-general-linear-recurrence-relations/ I know @mukushla and @joemath314159 are familiar with this technique

OpenStudy (anonymous):

@zscdragon what class is this for?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (1+5^n)\]

OpenStudy (anonymous):

i am also confused on how you would get the 20th term

OpenStudy (asnaseer):

ok, after reading about recurrence relations on the link I gave you above, this is what I found. we have: \(a_n=5a_{n-1}+1\) with \(a_0=0\) we first re-arrange this to: \(a_n-5a_{n-1}=1\) the solution for this type of recurrence relation is of the form: \(a_n=a_{n1}+a_{n2}\) where \(a_{n1}\) is the solution to: \(a_n-5a_{n-1}=0\) which will be of the form:\[a_{n1}=A5^n\] and \(a_{n2}\) is a polynomial solution of the form defined by the degree of the polynomial on the right-hand-side of \(a_n-5a_{n-1}=1\), so in this case we can say:\[a_{n2}=B\]where B is some constant. we then substitute \(a_{n2}\) into this equation to get:\[B-5B=1\implies B=-\frac{1}{4}\] putting these together we get:\[a_n=a_{n1}+a_{n2}=A5^n+B=A5^n-\frac{1}{4}\]and we know \(a_0=0\) therefore:\[0=A-\frac{1}{4}\implies A=\frac{1}{4}\]finally we get:\[a_n=\frac{5^n-1}{4}\]this can then be used to find the 20'th term. I hope I did this correctly. maybe @mukushla and @joemath314159 can check my work. more importantly, I hope you understood the process @zscdragon?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

why did you tag me @nickhouraney ?

OpenStudy (anonymous):

Looks good to me :)

OpenStudy (asnaseer):

@joemath314159 - thanks for confirming :)